Car Survey: In a survey of 19000 people who owned a certain type of​ car, 1045 said they would buy that type of car again. What percent of the people surveyed were satisfied with the​ car?
Answer by math_tutor2020(3816)


A = number of people who like the car
B = number total

Divide the two values.
A/B = 1045/19000 = 0.055

That converts to the percentage 5.5% when moving the decimal point two spots to the right.
This is equivalent to multiplying by 100.